# Difficult control is related to instability in biologically inspired Boolean networks

### v1

@@
Abstract Previous work in Boolean dynamical networks has suggested that the number of components that must be controlled to select an existing attractor is typically set by the number of attractors admitted by the dynamics, with no dependence on the size of the network. Here we study the rare cases of networks that defy this expectation, with attractors that require controlling most nodes. We find empirically that unstable fixed points are the primary recurring characteristic of networks that prove more difficult to control. We describe an efficient way to identify unstable fixed points and show that, in both existing biological models and ensembles of random dynamics, we can better explain the variance of control kernel sizes by incorporating the prevalence of unstable fixed points. In the end, the fact that these exceptions are associated with dynamics that are unstable to small perturbations hints that they are likely an artifact of using deterministic models. These exceptions are likely to be biologically irrelevant, supporting the generality of easy controllability in biological networks.
@@ 1 Introduction
Since Kauffman’s original proposal [1] that the attractors of network dynamics could effectively represent different cell types, Boolean networks have found broad application in theoretical biology as the most parsimonious mathematical representation of real biological networks [2, 3].
In Ref. [4], we demonstrated that dynamic Boolean networks are often easy to control. We illustrated this on a large database of experimentally derived biological networks, as well as ensembles of random networks produced through several generating mechanisms. While various approaches have been proposed for quantifying the controllability of complex networks [5, 6, 7, 8, 9], we focus on the difficulty of steering the dynamics towards the attractors of the unconstrained system [6], rather than navigating through all possible states [10]. This matches with the assumption that biological function is mostly captured by attractors, in that they represent the phenotypically and functionally distinct cell types expressed by the regulatory network [1].
@@
Our findings reveal that, contrary to intuition, control toward existing attractors does not inherently become more difficult as the size of the network increases. Instead, a significantly more crucial factor is the number of attractor states generated by the dynamics. We find that the control needed to select one of the attractors typically grows logarithmically with the number of attractors. This concept is made precise by defining a control kernel: a set of nodes of minimal size whose states can be pinned to fixed values to force the system into a given attractor, regardless of initial conditions [11]. Our result can then be summarized in the statement that the average size of the control kernel of a dynamical Boolean network, $\langle|\textrm{CK}|\rangle$, typically scales logarithmically with the number of its attractors, $r$:
@@
In Ref. [4], we reinterpreted this result in terms of the conjectured expectation value of the witness set in computational learning theory [12]. More importantly, we view it as the most significant theoretical justification to date for the feasibility of genetic reprogramming [13].
@@
While networks that were much more difficult to control than predicted by our scaling law were limited in number (roughly 2% of tested networks), we noted their occurrence in Ref. [4]. Specifically, some networks had significant numbers of attractors whose control required pinning nearly all nodes in the network. The presence of such networks, along with the unaccounted reasons for their difference from more easily controllable networks, could undermine one’s confidence in predicting the required level of control based solely on knowledge of a network’s attractors, as well as the hypothesis that a network’s size is typically irrelevant to the difficulty of control.
@@
In this work, we investigate the factors that contribute to the increased difficulty in controlling exceptional networks. Our goal is to find approximations to the mean control kernel size that are simple to compute and do not rely on solving the full dynamics, which becomes intractable in moderately sized networks.
@@
As a first step, we expand our database of biological networks by incorporating additional networks collected in Ref. [14], representing the most extensive collection of Boolean models for biological networks at the time of writing this manuscript. None of the new networks exhibit a significant deviation from our original predictions, leaving us with the task of explaining only the exceptions previously identified.
@@
A direct examination of these exceptions highlights the role played in the controllability of the network by the presence of isolated fixed points, i.e., fixed points of the dynamics that do not attract any other state. We establish a strong correlation between this characteristic and the difficulty of control, and we use this as the rationale for a correction to our original scaling law (Equation 1). The revised scaling law relies not only on the knowledge of the number of attractors but also on the number of isolated fixed points among them. Interestingly, the latter can be easily evaluated even for networks where not all basin sizes are known, particularly in cases where the identity of all attractors can be obtained by exploiting modularity in the network topology. We test whether our new empirical formula outperforms our previous scaling law, and demonstrate that it accounts for the majority of the remaining variance not captured by the original scaling.
@@ 2 Results
To explore the drivers of exceptionally large control kernels, we first ask: What is the maximum possible control kernel size for a given network? Given our definition of control, fixed-point attractors always have a control kernel, as control can be attained by pinning all nodes in the network. Note that cyclic attractors are sometimes not able to be forced with static control because we are limited to pinning non-cycling nodes — in these cases, we say that a control kernel does not exist. This sets the simplest upper bound on the size of control kernels for controllable attractors of $|\mathrm{CK}|\leq n$. A more economical approach avoids pinning nodes in any peripheral trees that are wholly dependent on nodes within closed feedback loops in the regulatory network topology. We will refer to the network’s core as the set of nodes that exclude any peripheral trees, with a number of core nodes $n_{\mathrm{c}}$. As control kernels are defined as sets of controlling nodes of minimal size, we can ensure a tighter bound of $|\mathrm{CK}|\leq n_{\mathrm{c}}$.
We might naively expect that it would be more difficult to steer the system toward attractors that are less likely to be found when initializing the system in a random state. Therefore, one might expect extreme control kernel sizes for attractors with very small basins of attraction (defined as the states that lead to each attractor under the dynamics). While basin sizes are not generally easy to compute, it is relatively simple to highlight the most extreme scenarios when an attractor is an “isolated” fixed point, i.e., with a basin size of 1 (see Methods).
@@
Both isolated fixed points and maximal control kernel sizes ($|\mathrm{CK}|=n_{\mathrm{c}}$) are connected to concepts of instability (Fig. 1). First, fixed points with maximal control kernels are locally unstable, in the sense that such a fixed point must be unstable to individual bit flips. That is, each state at Hamming distance 1 from the fixed point must be in the basin of a different attractor (otherwise one of the nodes could be left unpinned and the system would still return to the fixed point). Note that local instability does not necessarily imply that no other states transition into the fixed point under the dynamics. Also, while having a maximal control kernel implies local instability, local instability does not in itself guarantee a maximal control kernel. Second, isolated fixed points are globally unstable, in the sense that no other states transition into them. [^1] This is distinct from other notions of instability in the literature on Boolean networks, such as instability of cycles to delays in updating [15]. Though a fixed point being globally unstable does not directly imply that it has a maximal control kernel, we might expect some correlation between instability and difficulty of control.
@@
In Fig. 2, we compare average control kernel sizes in networks that have isolated (globally unstable) fixed points to the distribution over all networks (A and B) and control kernel sizes for individual isolated fixed points compared to the distribution over all attractors (C and D). The control kernel sizes of isolated fixed points are often close to $n_{\mathrm{c}}$, matching with our intuition. Some of the biological networks have no feedback loops and are therefore controlled entirely by their input nodes, which trivially creates isolated fixed points with $|\mathrm{CK}|=n_{\mathrm{c}}$ (hatched areas in Fig. 2). Even aside from this effect, isolated fixed points tend to have very large control kernels.
@@
Figure 3: A correction that incorporates the number of isolated fixed points leads to better predictions of control kernel sizes. (A and B) The simple prediction that the mean control kernel size within each network is equal to the logarithm of the number of attractors $r$ (Eq. 1) works well for most networks, but there are a number of exceptional networks that are much harder to control than expected. These exceptional networks all have isolated fixed points (highlighted with a darker color). (C and D) A correction that assumes isolated fixed points have control kernel of size $n_{\mathrm{c}}$ (Eq. 2) leads to better predictions.
@@
where $s$ is the number of isolated fixed-points. As shown in Fig. 3, this simple estimate does quite well in most cases, outperforming Eq. 1 for almost all networks we tested.
The significance of Eq. 2 lies in the fact that, similar to Eq. 1, it enables an estimation of the average control kernel size solely based on the characteristics of the attractor landscape of the dynamics. This is achieved without the necessity of solving the NP-hard problem of determining control kernels for all $r$ attractors.
@@
Significantly, Eq. 2 introduces a dependence of the control kernel on the network size (specifically on $n_{\mathrm{c}}$). This seems to be at odds with our findings in Ref. [4], where we concluded that the average control kernel size remains independent of the network size. It is crucial to note that the parameter $s$ is typically small in comparison to $r$, so that the influence of $n_{\mathrm{c}}$ is often hidden upon considering an average across all attractors in a network. This becomes especially apparent when Eq. 2 is recast in the form
@@ 3 Methods
Boolean network dynamics and control kernels. In a Boolean network, the states of individual components in a system are represented as Boolean variables that are updated in discrete timesteps according to fixed logical rules. Here we use a synchronous updating scheme, which creates deterministic dynamics. After transient dynamics, all initial states lead to a set of attractors that represent the possible long-term behaviors of the system. Each attractor consists either of a single state (fixed point) or cycles through a larger set of states.
@@
We characterize the difficulty of control using the notion of a control kernel [11]. Specifically, we define a control kernel for a given attractor as a set of nodes of minimal size that can be statically pinned to fixed values such that, in the pinned dynamics, all initial conditions converge to the desired attractor. Using this static definition of control, all fixed point attractors have a control kernel of some size, but not all cycles are controllable.
@@
For set environmental conditions (i.e. fixed states of input nodes [4]) multiple attractors can only exist if the network structure presents feedback loops [6]. Any set of nodes connected in a tree structure without closed loops will depend deterministically on the states of input nodes in each attractor state. For this reason, nodes outside of a feedback loop will never be part of a control kernel (as all behavior of trees will be fully controlled once the attractor is selected within the core). For this reason, we remove from our analysis all dependent trees of nodes from each network, leaving only a number $n_{\mathrm{c}}\leq n$ of “core” nodes that participate in feedback loops and have the potential to be part of control kernels. Alternatively, we could include all $n$ nodes in our analysis, but this would preclude the possibility of having any isolated fixed points in networks that have any nodes on which no other nodes depend (as changing the states of such nodes does not affect the dynamics).
@@
To increase the variance in our tested models, we also include three ensembles of random Boolean networks, as presented in Ref. [4]. These include Kauffman networks with random truth tables, as well as Erdos–Renyi networks governed by threshold rules, with thresholds chosen either at zero total input or balanced to account for the number of inputs to each node (for details, see Ref. [4]).
@@
Let us refer to a generic Boolean network with $n$ nodes and updating rules
$x^{\prime}_{1}$ $=$ $f_{1}(x_{1},\dots,x_{n})$
$\dots$
$x^{\prime}_{n}$ $=$ $f_{n}(x_{1},\dots,x_{n})\ ,$ (4)
where $x_{i}$ is the value (0 or 1) of node $i$ at time $t$, and $x^{\prime}_{i}$ is its value at time $t+1$. We will call $S_{i,b}$ the subset of the configuration space $S$ containing all the states satisfying the non-linear equation $f_{i}(x_{1},\dots,x_{n})=b$, where $b=0,1$. Therefore, given a state $X=(x_{1},\dots,x_{n})$, its backward reachable set, defined as the set of its pre-images, is
@@
$P_{X}=P_{1,x_{1}}\bar{\times}P_{2,x_{2}}\bar{\times}\dots\bar{\times}P_{n,x_{n}}\ ,$

### v2

@@
Abstract Previous work in Boolean dynamical networks has suggested that the number of components that must be controlled to select an existing attractor is typically set by the number of attractors admitted by the dynamics, with no dependence on the size of the network. Here we study the rare cases of networks that defy this expectation, with attractors that require controlling most nodes. We find empirically that unstable fixed points are the primary recurring characteristic of networks that prove more difficult to control. We describe an efficient way to identify unstable fixed points and show that, in both existing biological models and ensembles of random dynamics, we can better explain the variance of control kernel sizes by incorporating the prevalence of unstable fixed points. In the end, the association of these outliers with dynamics that are unstable to small perturbations reveals them as artifacts of deterministic models, making them less biologically relevant and reinforcing the generality of easy controllability in biological networks.
@@ 1 Introduction
The formulation of an exhaustive theory for controlling complex biological systems, and gene regulatory networks (GRNs) more specifically, remains a key objective in systems biology [1, 2]. And while optimal control theory offers a broad set of tools for various applications in mathematics and engineering, its traditional assumptions often require modification for biological contexts. Biological processes are typically inherently nonlinear [3], and the focus is frequently restricted to coarse-grained output states [4], rather than driving the system to an arbitrary state [5]. There are a number of approaches to quantifying control in biology, which vary in their basic assumptions. States can be modeled as continuous or discrete, dynamics can be assumed to be stochastic or deterministic, control inputs can be static or dynamic, and the goal of control can be to reach arbitrary states or to reach particular phenotypes (attractors). Much work in biological control has focused on phenotypic control toward specific coarse-grained output states [4], as it is arguably more relevant to biological function than driving a system into any arbitrary state [5]. Some work inspired by traditional control theory has focused on control toward arbitrary states, either in continuous [6] or discrete state space [7, 8], though this typically leads to control strategies that are significantly more difficult to implement than what is necessary to select a single desired output. The simplicity of discrete state spaces (and the inability to constrain parameters in continuous formulations) has guided many to take a Boolean approach. The typical interventions available in experiments has led to a focus on control inputs that are largely static, either applied permanently as in a gene knockout or temporarily over some timescale [9, 10, 11, 12, 13], though some approaches allow for arbitrary dynamics of inputs [7, 11]. Finally, for simplicity and mathematical tractability, the updating functions of Boolean network models are often defined deterministically. A mild form of stochasticity is sometimes incorporated in defining the order in which nodes are updated across time (e.g., [9, 10, 11, 13]), or noise can be modeled in the form of bit-flips or probabilistic transitions [14].
We focus here on discrete, Boolean, deterministic, static, phenotypic control. In this context, an important concept for regulating network dynamics is the control kernel (CK). First proposed in [15], the CK refers to a set of genes of minimal size whose external manipulation is enough to guide the network towards a target steady-state gene activation pattern, i.e. an attractor of the dynamical model of the GRN. The CK is defined such that it can force a particular attractor state regardless of the initial condition. This matches with the assumption that biological function is mostly captured by attractors, in that they represent phenotypically and functionally distinct cell types [16].
Several related approaches have been proposed for the study of the optimal control of deterministic Boolean networks. Stable motif analysis defines control sets by detecting positive circuits within the network that can sustain trap spaces in the dynamics [17]. Feedback vertex sets offer an efficient upper bound on the size of minimal control sets in a way that does not depend on the specific dynamics governing each node [18, 19, 5].
A simple heuristic argument suggests that the size of CKs may scale approximately with the logarithm of the number of original attractors. Under the simplifying assumption that state transitions are all equally likely, forcing a single node into a specific state would generally halve the number of attainable states. Therefore, by controlling $c$ nodes, we would expect to reduce $2^{c}$ possible attractors down to a single attractor, suggesting the logarithmic scaling of CK size. However, the task of identifying the minimal set of controlling nodes is nontrivial and has been proven NP-hard by Akutsu et al. [20], and counterexamples can be constructed that require much larger CKs.
@@
In Ref. [21], we demonstrated that dynamic Boolean networks do often have CK sizes that scale logarithmically with the number of attractors. We illustrated this on a large database of experimentally derived biological networks, as well as ensembles of random networks. These previous findings revealed that control toward existing attractors does not inherently become more difficult as the size of the network increases. Instead, the number of attractor states generated by the dynamics is a much better predictor of the difficulty of control than the network size. Our result could then be summarized in the statement that the average size of the control kernel of a dynamical Boolean network, $\langle|\textrm{CK}|\rangle$, typically scales logarithmically with the number of its attractors, $r$:
@@
In Ref. [21], we reinterpreted this result in terms of the conjectured expectation value of the witness set in computational learning theory [22]. More importantly, we view it as one of the most significant theoretical and empirical justifications to date for the feasibility of genetic reprogramming [23].
Unbeknownst to us, research conducted by Akutsu’s group and published earlier than our work had reached a very similar conclusion and independently identified the logarithmic scaling described, even if under somewhat different assumptions and on more abstract models [24]. A number of clever simplifications make the mathematics more manageable and allow for the direct derivation of logarithmic scaling as in Eq. 1. The primary difference between the control defined by Hou et al. and our study is the duration of the pinning. Hou et al. assume that the pinning, which constitutes the control signal, lasts for a single timestep and is then removed, whereas we assume that the pinning is permanent. Our choice of permanent pinning is influenced by our focus on biological networks, where long-timescale perturbations, such as those observed in gene overexpression, knockdown, and knockout experiments, have practical significance. Another important difference is that Hou et al. assume that the basins of all attractors are of equal size. Due to the large variability in the basin sizes of the experimentally motivated models we examine, our current work aims to shed more light on the still hazy relationship between basin size and controllability.
@@
In our previous study, some networks were much more difficult to control than predicted by a logarithmic scaling law (roughly 2% of tested networks). A few networks had significant numbers of attractors whose control required pinning nearly all nodes in the network. The presence of such networks, along with the unaccounted reasons for their difference from more easily controllable networks, could undermine one’s confidence in predicting the required level of control based solely on knowledge of a network’s attractors, as well as the hypothesis that a network’s size is typically irrelevant to the difficulty of control.
@@
Our focus in this work is to investigate the factors that contribute to the increased difficulty of controlling such networks. Our goal is to find approximations to the mean control kernel size that are simple to compute and do not rely on solving the full dynamics, which becomes intractable in moderately sized networks.
@@
As a first step, we expand our database of biological networks by incorporating additional networks collected in Ref. [25], representing the most extensive collection of Boolean models for biological networks at the time of writing this manuscript. None of the new networks exhibit a significant deviation from our original predictions, leaving us with the task of explaining only the outliers previously identified.
@@
A direct examination of these outlier networks highlights the presence of isolated fixed points, i.e., fixed points of the dynamics that do not attract any other state. We establish a strong correlation between this characteristic and the difficulty of control, and we use this as the rationale for a correction to our original scaling law (Equation 1). The revised scaling law relies not only on the knowledge of the number of attractors but also on the number of isolated fixed points among them. Interestingly, the latter can be easily evaluated even for networks where not all basin sizes are known, particularly in cases where the identity of all attractors can be obtained by exploiting modularity in the network topology. We test whether our new empirical formula outperforms our previous scaling law, and demonstrate that it accounts for the majority of the remaining variance not captured by the original scaling.
@@ 2 Results
To explore the drivers of unusually large control kernels, we first ask: What is the maximum possible control kernel size for a given network? Given our definition of control, fixed-point attractors always have a control kernel, as control can be attained by pinning all nodes in the network. Note that cyclic attractors are sometimes not able to be forced with static control because we are limited to pinning non-cycling nodes — in these cases, we say that a control kernel does not exist. This sets the simplest upper bound on the size of control kernels for controllable attractors of $|\mathrm{CK}|\leq n$. A more economical approach avoids pinning nodes in any peripheral trees that are wholly dependent on nodes within closed feedback loops in the regulatory network topology. We will refer to the network’s core as the set of nodes that exclude any peripheral trees, with a number of core nodes $n_{\mathrm{c}}$. As control kernels are defined as sets of controlling nodes of minimal size, we can ensure a tighter bound of $|\mathrm{CK}|\leq n_{\mathrm{c}}$.
We might naively expect that it would be more difficult to steer the system toward attractors with a smaller than average basin size when initializing the system in a random state. Therefore, one might expect extreme control kernel sizes for attractors with very small basins of attraction (defined as the states that lead to each attractor under the dynamics). While basin sizes are not generally easy to compute, it is relatively simple to highlight the most extreme scenarios when an attractor is an “isolated” fixed point, i.e., with a basin size of 1 (see Methods).
Figure 1: Examples of networks with and without exceptional states. Networks that have control kernels of maximal size $n_{c}$ (“Max CK”) are related to those that have locally and globally unstable states. Empirically, most networks do not have control kernels of maximal size (first row). All observed networks that we identify as being unusually difficult to control contain both fixed points with maximal CK size and globally unstable fixed points (second row). Some of the biological networks have no cycles in their interaction networks, which trivially creates $2^{m}$ fixed points with $|\mathrm{CK}|=m=n_{c}$; since they follow Eq. 1, we do not consider these networks to be unusually difficult to control (third row). Finally, it is possible to construct networks that contain fixed points with maximal CK size but no globally unstable fixed points, but we find no examples of this case in the sampled networks (fourth row). Globally unstable fixed points have an incoming edge only from themselves in the state transition graph. Locally unstable fixed points can have other incoming edges, but none from nearest Hamming neighbors (gray region). Fixed points with maximal CK size must be locally unstable.
@@
Both isolated fixed points and maximal control kernel sizes ($|\mathrm{CK}|=n_{\mathrm{c}}$) are connected to concepts of instability (examples shown in Fig. 1 and summarized in Fig. 2). First, fixed points with maximal control kernels are locally unstable, in the sense that such a fixed point must be unstable to individual bit flips. That is, each state at Hamming distance 1 from the fixed point must be in the basin of a different attractor (otherwise one of the nodes could be left unpinned and the system would still return to the fixed point). Note that local instability does not necessarily imply that no other states transition into the fixed point under the dynamics (a counterexample is shown in the last row of Fig. 1). Also, while having a maximal control kernel implies local instability, local instability does not in itself guarantee a maximal control kernel. Second, isolated fixed points are globally unstable, in the sense that no other states transition into them. [^1] This is distinct from other notions of instability in the literature on Boolean networks, such as instability of cycles to delays in updating [26]. Though a fixed point being globally unstable does not directly imply that it has a maximal control kernel, we might expect some correlation between instability and difficulty of control.
@@
In Fig. 3, we compare average control kernel sizes in networks that have isolated (globally unstable) fixed points to the distribution over all networks (A and B) and control kernel sizes for individual isolated fixed points compared to the distribution over all attractors (C and D). The control kernel sizes of isolated fixed points are often close to $n_{\mathrm{c}}$, matching with our intuition. Some of the biological networks have no feedback loops and are therefore controlled entirely by their input nodes, which trivially creates isolated fixed points with $|\mathrm{CK}|=n_{\mathrm{c}}$ (hatched areas in Fig. 3 and third row in Fig. 1). Even aside from this effect, isolated fixed points tend to have very large control kernels.
@@
Figure 4: A correction that incorporates the number of isolated fixed points leads to better predictions of control kernel sizes. (A and B) The simple prediction that the mean control kernel size within each network is equal to the logarithm of the number of attractors $r$ (Eq. 1) works well for most networks, but there are a number of outlier networks that are much harder to control than expected. We mark networks as outliers (triangles) when their mean control kernel size is more than $3\sigma$ away from the expected value of $\log_{2}{r}$, where $\sigma$ is the standard deviation of the residuals $\langle|\textrm{CK}|\rangle-\log_{2}{r}$. These outlier networks all have isolated fixed points (highlighted with a darker color). (C and D) A correction that assumes isolated fixed points have control kernel of size $n_{\mathrm{c}}$ (Eq. 2) leads to better predictions. For the biological networks, the RMSE is 0.93 when using Eq. 1 and reduces to 0.60 when using Eq. 2. For the random networks, the RMSE reduces from 1.3 to 0.68.
@@
where $s$ is the number of isolated fixed-points. As shown in Fig. 4, this simple estimate does quite well in most cases, outperforming Eq. 1 for almost all networks we tested. In particular, if we define networks that are unusually difficult to control as those that are outliers in Fig. 4A and B (shown as triangles, and defined as those with residuals $|\langle|\textrm{CK}|\rangle-\log_{2}{r}|>3\sigma$, with $\sigma$ the standard deviation of those residuals across networks in each ensemble), Eq. 2 successfully corrects the predictions of mean control kernel sizes for those difficult cases.
The significance of Eq. 2 lies in the fact that, similar to Eq. 1, it enables an estimation of the average control kernel size solely based on the characteristics of the attractor landscape of the dynamics. This is achieved without the necessity of determining control kernels for all $r$ attractors.
@@
Significantly, Eq. 2 introduces a dependence of the control kernel on the network size (specifically on $n_{\mathrm{c}}$). This seems to be at odds with our findings in Ref. [21], where we concluded that the average control kernel size remains independent of the network size for a given $r$. It is crucial to note that the parameter $s$ is typically small in comparison to $r$, so that the influence of $n_{\mathrm{c}}$ is often hidden upon considering an average across all attractors in a network. This becomes especially apparent when Eq. 2 is recast in the form
@@
Finally, treating Eqs. 1 and 2 as statistical predictors of average control kernel size, we examine the distributions of residuals away from these predictions. The non-normality of the residual distribution for the uncorrected Eq. 1 is visible in Fig. 5B, with the outlier networks contributing to a sizable skew. This gives a statistical motivation to look for a better predictor. Our corrected Eq. 2 then produces a distribution of residuals that is closer to normal, with the outliers largely gone (Fig. 5C). We quantify this approach to normality using the Shapiro-Wilk statistic: for the biological networks, $W$ increases from $0.64$ using Eq. 1 to $0.85$ using Eq. 2, and for the random networks, $W$ increases from $0.68$ to $0.87$. See SI section S3 and Fig. S2 for a more thorough discussion and visualization of the residual distributions.
@@ 3 Methods
Boolean Networks and attractor dynamics. In this section, we define key terms and concepts related to Boolean dynamical systems. Readers already familiar with Boolean network dynamics may skip to the next section.
The state of a Boolean network at time $t$ is characterized by the states of its $n$ nodes, $x_{i}(t)$, where $i=1,\dots,n$. Due to the Boolean nature of the model, each node $i$ can be in one of two possible states: ON, represented by $x_{i}(t)=1$, or OFF, represented by $x_{i}(t)=0$. In the context of cell regulation, the nodes typically correspond to interacting genes, with $x_{i}(t)$ serving as the Boolean approximation of the expression level of gene $i$. Here, the gene is considered expressed when $x_{i}(t)=1$, and inactive otherwise.
The dynamics of a deterministic Boolean network are characterized by $n$ Boolean functions $f_{1},\dots,f_{n}$, which take as input the array $x_{1}(t),\dots,x_{n}(t)$ and produce the network’s configuration at time $t+1$:
$x_{1}(t+1)$ $=$ $f_{1}(x_{1}(t),\dots,x_{n}(t))$
[…]
$x_{n}(t+1)$ $=$ $f_{n}(x_{1}(t),\dots,x_{n}(t))\ .$ (4)
The specific form of these Boolean functions can be identified through experiments [27, 28]. This setup represents a synchronous update of all nodes in the network.
With $n$ nodes, there are $2^{n}$ potential network configurations (i.e., gene expression patterns in the context of a genetic network). We will refer to these configurations as states of the network and the collection of all possible network states as the configuration space. The dynamics of the network are depicted by a time series of states. Although large, the configuration space is finite. Hence, when the functions $f_{i}$ are deterministic, these dynamical trajectories will eventually settle into either a fixed state or a cycle of states. These specific sets of states are termed the attractors of the Boolean network. The collection of initial states that converge to a particular attractor is called its basin of attraction.
@@
Control kernels. We characterize the difficulty of control using the notion of a control kernel [15]. Specifically, we define a control kernel for a given attractor as a set of nodes of minimal size that can be statically pinned to fixed values such that, in the pinned dynamics, all initial conditions converge to the desired attractor. Using this static definition of control, all fixed point attractors have a control kernel of some size, but not all cycles are controllable.
Many biological network models include nodes that represent external or environmental conditions. We call such nodes “input nodes,” and we implement their dynamics using the updating rule $x(t+1)=x(t)$. Input nodes are included in the network state, and we incorporate all possible values of these nodes when enumerating a network’s attractors. Based on our definition of control, all input nodes are included in every control kernel, as no other nodes can influence their state [21]. In rare cases, a network model will define a node’s dynamics such that it always results in a specific constant value $C$, that is, $x(t+1)=C$, with $C=0$ or $1$. Such nodes have a fixed value across all possible attractors. They therefore do not distinguish different attractors and are thus never included as part of a control kernel.
@@
For set environmental conditions (fixed states of input nodes) multiple attractors can only exist if the network structure presents feedback loops [29]. Any set of nodes connected in a tree structure without closed loops will depend deterministically on the states of input nodes in each attractor state. For this reason, nodes outside of a feedback loop will never be part of a control kernel (as all behavior of trees will be fully controlled once the attractor is selected within the core). For this reason, we remove from our analysis all dependent trees of nodes from each network, leaving only a number $n_{\mathrm{c}}\leq n$ of “core” nodes that participate in feedback loops and have the potential to be part of control kernels.222 A subtlety arises in computing $n_{c}$ in that redundant expressions in the updating rules can create inoperative edges [30]. We detail in supplementary section S2 why we do not expect this to affect our results. Alternatively, we could include all $n$ nodes in our analysis, but this would preclude the possibility of having any isolated fixed points in networks that have any nodes on which no other nodes depend (as changing the states of such nodes does not affect the dynamics).
Finally, we note that the core nodes of a network with no cycles consist only of the network’s input nodes. This corresponds to the fact that every attractor is fully specified by the states of the inputs, and results in the control kernel being equal to the set of input nodes. It is also possible for inputs to fully control networks that include cycles. In all of these cases, the number of attractors is $2^{m}$, control kernels are of size $m$, and the logarithmic scaling of Eq. (1) is satisfied exactly. Therefore, these cases do not show up among those networks that we consider unusually difficult to control.
@@
A key limitation of Boolean dynamical models is the need to limit their size, as the number of network states grows exponentially with the number of nodes. To manage this, peripheral regions of the interactome are often simplified into input and output nodes, based on the assumption that the highlighted subsystems function relatively independently of the complete cellular interactome. While studies have shown that the network topology of these reduced models partially preserves the motif distribution of the full networks [32], unsurprisingly this preservation diminishes as the size of the extracted subgraphs decreases.
To increase the variance in our tested models, we also include three ensembles of random Boolean networks, as presented in Ref. [21]. These include networks with random truth tables, as well as Erdös-Rényi networks governed by threshold rules, with thresholds chosen either at zero total input or balanced to account for the number of inputs to each node. First, we use the well-studied ensemble of p-K random networks [33] to generate dynamics with random truth tables. Each node receives input from $K=2$ nodes, with Boolean updating rules chosen such that the probability of the ON state is $p$. We sampled 225 networks from this ensemble, with 75 each having $p=0.25$, $0.5$, and $0.75$, and 25 each within these sets having number of nodes $n=10$, $15$, and $20$. We found control kernels for all controllable attractors in each of these sampled networks. Second, we define an ensemble that assigns each node a set of input nodes whose states are summed and compared to a threshold [34, 15]. The network’s dependency structure $A$ is chosen as an Erdös-Rényi graph with average degree $d$, and each edge in this graph is assigned with probability $p_{I}$ a value of $-1$ (representing an inhibitory interaction) or $+1$ (representing an excitatory interaction). Each node’s state is determined by comparing the sum of the incoming signed inputs $s_{i}=\sum_{j}A_{ij}x_{j}(t)$ to its threshold $\tau_{i}$:
$x_{i}(t+1)=\begin{cases}0,\textrm{if}~{}s_{i}<\tau_{i}\\ x_{i}(t),\textrm{if}~{}s_{i}=\tau_{i}\\ 1,\textrm{if}~{}s_{i}>\tau_{i}.\end{cases}$ (5)
For one ensemble, we set thresholds to zero ($\tau_{i}=0~{}\forall~{}i$). In cases with little inhibition, these networks are biased toward excitation. A second ensemble consists of “balanced” threshold networks, with thresholds of individual nodes set at the center of the distribution of possible summed inputs ($\tau_{i}=\frac{1}{2}\sum_{j}A_{ij}$). We sampled 75 networks using each type of threshold, with each combination of $d=\{1,2,3,4,5\}$, $p_{I}=\{0.1,0.3,0.5,0.7,0.9\}$, and $n=\{10,15,20\}$, for a total of 150 networks. Of these, for 146 we were successful in finding control kernels for all controllable attractors.
The control results for all random networks included in this study are shown in Fig. S1, where we replot Fig. 4B to highlight the differences arising from the various methods used to construct the random networks. Notably, all outliers, marked as triangles in Fig. 4B, originate from the zero-threshold ensemble.
@@
Let us refer to a generic Boolean network with $n$ nodes and updating rules as in (4). We will call $S_{i,b}$ the subset of the configuration space $S$ containing all the states satisfying the non-linear equation $f_{i}(x_{1},\dots,x_{n})=b$, where $b=0,1$. Therefore, given a state $X=(x_{1},\dots,x_{n})$, its backward reachable set, defined as the set of its pre-images, is
@@
$P_{X}=P_{1,x_{1}}\bar{\times}P_{2,x_{2}}\bar{\times}\dots\bar{\times}P_{n,x_{n}}\ ,$ (7)